# New Kid in the Classroom: Exploring Student Perceptions of AI Coding Assistants

### v3

@@ 4 Discussion
Overall, these results underscore the need for thoughtful AI integration in programming education. Scaffolding activities that require students to adapt or critique AI-generated code may help mitigate overreliance and support the development of core skills. While this study is limited by its small sample size and specific context, the findings present a compelling case for future research with a focus on longitudinal data and objective performance metrics to better understand the long-term pedagogical impact of AI-assisted learning.
@@ 5 Conclusions
As generative AI tools become increasingly embedded in programming education, understanding their impact on learning is both urgent and complex. This exploratory study offers preliminary evidence that, while students find AI support valuable and report increased confidence during assisted tasks, these perceptions do not always align with independent performance or comprehension. Discrepancies between self-reports and instructor observations suggest that learners may underestimate their reliance on AI or overestimate their own mastery.
Such findings underscore the importance of designing AI-integrated activities that not only assist but also challenge students to internalize concepts and apply them independently. Educational strategies should aim to leverage AI for scaffolding, while encouraging active engagement and critical thinking. Mixed-method approaches—combining survey, behavioral, and performance data—will be essential to evaluate the true pedagogical impact of AI assistance.
Moreover, incorporating AI into the programming classroom is not simply a matter of access or efficiency, but of pedagogy. The goal should not be to replace effort with automation, but to use automation wisely—to support, not supplant, the learning process. Blending AI assistance with opportunities for independent problem-solving may help strike this balance, offering a promising direction for future research. In this new era of AI coding tools, it is essential to rethink programming education through strategies that foster engagement and gradually shift responsibility to the learner. Ultimately, the student perceptions captured in this study present a compelling case for reimagining paths that facilitate acquiring and developing computational thinking skills. We fully support this new direction, particularly as AI coding assistants become increasingly prevalent.

### v4

@@ 4 Discussion
Overall, these results underscore the need for thoughtful AI integration in programming education. Scaffolding activities that require students to adapt or critique AI-generated code may help mitigate overreliance and support the development of core skills. We remark that despite being limited by its small sample size and specific context, this study’s findings provide a foundation for future research that incorporates behavioral, longitudinal, and performance metrics to better understand the long-term impact of AI-assisted learning.
@@ 5 Conclusions
The rapid adoption of generative AI in programming education raises critical questions about its long-term impact on learning. In this context, this exploratory study suggests that while students perceive AI support as valuable, their independent performance may not align with their sense of mastery, indicating a possible overestimation of their own skills. This disconnect highlights the potential for cognitive offloading, where reliance on the tool replaces the internalization of knowledge. These findings underscore the importance of designing activities that use AI for scaffolding, not to replace the essential practice of independent problem-solving. Ultimately, our results present a compelling case for reimagining programming education to foster active engagement and strengthen computational thinking.